OpenStudy (calculusxy):

Two books combined is $45 more than a desk. If the two books and desk is $145, what is the cost of one book. @ganeshie8

OpenStudy (campbell_st):

well let the dexk cost x, then the 2 books cost x + 45 so you have x + x + 45 = 145 solve for x then since the books cost x + 45 substitute you answer for x and then divide the total by 2 to get the cost of 1 book

OpenStudy (calculusxy):

Can you please explain it to me more? I don't understand this too much. @campbell_st

OpenStudy (calculusxy):

@jdoe0001 can you plz help?

OpenStudy (jdoe0001):

hmm.. looking

OpenStudy (jdoe0001):

well, I see it this way since both books seem to cost the same amount b = book d = desk so b + b = 2 books 2 books combined, cost whatever the desk costs plus $45 \(\bf b + b = d + 45\) if the 2 books and desk all together cost $145 \(\bf b + b + d = 145\) how much is 1 book? \(\bf b + b = d + 45 \implies 2b = d+45 \implies 2b-45 = d\\\quad \\ b + b + d = 145 \implies 2b +d = 145 \implies d = 145-2b\\ \quad \\ \quad \\ 2b-45 = d\qquad and\qquad d = 145-2b\\\quad \\ 2b-45 = 145-2b\)

OpenStudy (jdoe0001):

if you solve for "b", that'd be the price of a book

OpenStudy (calculusxy):

Can you help to solve for because I am not good at factoring.

OpenStudy (calculusxy):

for b I meant

OpenStudy (jdoe0001):

hmmmm, solving for "b" is just plain simplifying the exercise assumes you should know how to simplify it

OpenStudy (calculusxy):

thx

OpenStudy (jdoe0001):

yw

OpenStudy (calculusxy):

can u help me with one more?

OpenStudy (calculusxy):

@jdoe0001

OpenStudy (jdoe0001):

I can try

OpenStudy (calculusxy):

I will post a new one and mention u. thx